OpenStudy (anonymous):

Suppose you have two triangles. In triangleABC, AB = 4, BC = 5, and CA = 6. In triangleEFG, EF = 400, FG = 500, and GE = 600. Are these triangles similar, and if so, what is the scale factor? A. similar, scale factor = 10 B. similar, scale factor = 100 C. similar, scale factor = 1000 D. not similar

OpenStudy (anonymous):

Yes,\[ \frac{400}4=\frac{500}5=\frac{600}6=? \]

OpenStudy (anonymous):

\[ \frac{400}4=\frac{500}5=\frac{600}6=100 \]

OpenStudy (anonymous):

B is your answer